OpenStudy (anonymous):

Solve each proportion 5/ x+2=3/x+1

OpenStudy (campbell_st):

cross multiply which will give 5(x + 1) = 3(x + 2) distribute, collect like terms then solve for x

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

So what's the answer

OpenStudy (anonymous):

5/x+2=3/x+1 or.5(x+1)=3(x+2) or.5x+5=3x+6 or,5x-3x=6-5 or,2x=1 so,x=1/2 ans:

OpenStudy (anonymous):

Thank youn

OpenStudy (anonymous):

You